a block of weight 100 Newton`s is lying on a rough horizontal surface is coefficient of friction is one by root 3 .the least possible force that can move the block is

Grade:12th pass

1 Answers

Mahima Kanawat
6 years ago
The least possible force required is equal to the opposing force i.e., friction.Friction = umg= 1/√3× 100 = 100/√3 N ( WEIGHT = mg)